# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### v2

@@ 2.2.2 Retrieval-based Augmentation
Large Language Models have occasionally been observed to generate hallucinated or factually inaccurate text Zhou et al. (2021); Zhao et al. (2023). To mitigate these issues, several works have proposed to strengthen LLMs with information retrieval systems Shi et al. (2023); Jiang et al. (2023); Nakano et al. (2022), which enables LLMs to retrieve relevant contents from an external repository (knowledge corpus). It has been verified that retrieval-augmented LLMs can generate texts in response to user input with fewer hallucinations Nakano et al. (2022). Furthermore, by incorporating customized private data resources, retrieval-augmented LLMs can respond to in-domain queries that cannot be answered by LLMs trained with public data. As previous works Nakano et al. (2022); Shi et al. (2023); Jiang et al. (2023) suggest, LLMs can generate more factual answers by feeding the references retrieved from the external corpus. In TAP4LLM, we have fortified the document retrieval capabilities of LLMs and consider two components: (i) document references: to provide supplemental relevant web pages as the references for the given table; (ii) term explanation: to explain strange/ambiguous term in the given table. We utilize technologies like vector databases Wang et al. (2021a) and LangChain LangChain (2022) to facilitate the retrieval of pertinent information from Wikipedia [^1] https://www.wikipedia.org/. The details for document references and term explanation can be found in Appendix B.
@@ 2.2.3 Self-consistency-based Augmentation
We follow Sui et al. (2023) to provide the self-consistency-based augmentation approach. First, we append the instruction “Identify critical values and ranges of the last table related to the statement” to the initial prompt, and then forward this prompt to the LLM. The output generated from this instruction is then incorporated back into the prompt. Following this, we reintroduce the enriched prompt, now containing both the initial query and the newly generated insights, to the LLM along with the task-specific instructions for further processing.
@@ 2.3 Table Packing
The desire to maintain efficient reasoning without changing the LLMs architecture motivates us to propose the token allocations module. The packing component supports token-limit allocation for table sampling and augmentation. We conduct an empirical study to determine the proper proportion of the sub-table length and augmentation information length, as shown in Figure 3. The packing process is controlled by a user-defined parameter token limit, which determines the maximum truncate token length. Moreover, the study Sui et al. (2023) emphasizes a noteworthy observation regarding using markup languages like HTML or XML leads to much better generation quality by LLMs over TQA and TFV. In this pattern, TAP4LLM support multiple serialization functions, e.g., HTML, XML, JSON, CSV, NL+Sep (one of the typical options, e.g., using ‘||||’ as cell/column separator) and Markdown,etc.
3 Experiments
In this section, we first introduce the experiment settings, then we conduct extensive comparison between baseline models and each module in TAP4LLM. We further provide an ablation study of our method and a comprehensive evaluation on the usage of TAP4LLM. Please refer to Appendix DE for additional settings and experiments.
3.1 Experiment Settings
Dataset. We evaluate TAP4LLM on five TQA & TFV datasets: Sequential Question Answering (SQA) Iyyer et al. (2017), HybridQA Chen et al. (2020b), TabFact Chen et al. (2020a), ToTTo Parikh et al. (2020). we also set up TAP4LLM on a Text-to-SQL dataset Spider Yu et al. (2018), detailed in  E.4. The statistic of the datasets are given in Table 5, and the details of the datasets and metrics are described in Appendix D.1.
Models. We select state-of-the-art LLMs that have been widely studied in text generation and reasoning. Specifically, We focus on multiple GPT models and most updated open-source LLMs (Llama2-70B and Mixtral-8x7B) to test TAP4LLM effectiveness. The details for the tested models and the embedding methods can be found in Appendix D.2. The experiment results of open-sourced models can be found in Appendix E.3.
Table 2: Comparative results of the table augmentation methods. We use semantic-based sampling method without augmentation as the baseline. The term “Delta” refers to the performance gap between each method and the baseline. The top-3 performance gap on each dataset are highlighted in green, with the best performance being underlined. Noted that since only the ToTTo dataset contains hierarchical headers, we only provide the “header hierarchy” method on this dataset. “D/M + SF" refers to Dimension/Measure+ Semantic Field Type.
@@ 3.1 Experiment Settings
3.2 Comparison Results of Table Sampling
@@
According to Table 1, we conduct the comparative experiments on multiple table sampling methods and make several observations as follow: (1) Semantic-based sampling with column grounding outperforms other sampling methods across all datasets by effectively selecting table parts most relevant to queries. Centroid-based sampling also shows competitive results by clustering data points within tables, though it lacks query-table relevance consideration. Moreover, when combining these two strong variants together (hybrid sampling), it shows the most powerful capability. (2) The rule-based sampling method “content snapshot”, while not as precise in capturing query-specific information, offers a promising, efficient alternative by focusing on essential table content through $n$-gram overlap, without the need for complex embedding calculations. (3) In contrast, direct encoding methods, including using GPT-3.5-turbo with a 32k token limit or a 4K token-sized context window with truncation, demonstrate inferior performance. This suggests that while they can encompass more table information, they may introduce noise or lose critical context, undermining the table reasoning process and highlighting the importance of strategic sampling for optimal LLM performance.
3.3 Comparison Results of Augmentation.
@@
For the comparative experiments of table augmentation methods, we use the semantic-based sampling method as the baseline and report the performance gap between adding each augmentation method or not. According to Table 2, several insights can be found as follows: (1) Table augmentation methods further improve LLM’s reasoning ability after sampling. For example, “D/M + SF" achieves higher accuracy across all six datasets (most significant increase on TabFact +2.87%).“Docs References" and “Term Explanations" add meaningful context and semantic understanding to the model’s processing of tables, with (SQA +5.13%, ToTTo +4.07% ). The “self-prompting" further exemlifies the potential for iterative improvment in query and response generation. However, not all augmentation methods yield positive outcomes. “Table Size" offers minimal performance enhancement and “Header Hierarchy" shows that introducing a hierarchy may complicate the model’s ability to process the tabular information in some contexts, possibly by adding unnecessary complexity. (2) Additionally, the comparison of cell selection methods for “Term Explanations" highlights the superior performance of LLM-based selection over heuristic approaches. We find that LLM-based cell selection outperforms the heuristics-based cell selection with improvements in “Delta” ranging from 0.80% to 4.07%. While achieving higher performance, the LLM-based method also increases the calling budget as it requires additional LLM calls. These results indicate that the method’s effectiveness varies with the dataset. i.e. It’s beneficial for datasets requiring complex text understanding and generation (SQA and ToTTo). However, its impact is less distinct or even slightly negative in datasets involving different types of data or nuanced tasks (FEVEROUS and HybridQA).
@@
Through the experiment results, we also observe that different augmentation methods perform well on the same dataset. For example, “D/M + SF", “Statistics Feature", “Term Explanation" and “Self Prompting" all show significant improvement on the TabFact dataset. This suggests superposable effects through combining multiple augmentation methods may bring better performance. We only report the simplest way to append all the augmentation information together into the prompt, and leave the fine-grained way of combination for future investigation.
@@ 3.4 Ablation Study of TAP4LLM
As indicated in Table 3, we conduct an ablation study to assess the impact of various components on the performance of TAP4LLM. Each row represents the method performance without a certain component. We find that each component contributes to the model’s effectiveness. The study demonstrates that certain components, such as table sampling and table augmentation, are more crucial. Each dataset reacts differently to the removal of features, which highlights the necessity of a customized design when optimizing for particular datasets. We also report the performance using the most suitable combination of table sampling and augmentation for each dataset in Table 3.
3.5 Trade-offs between Token Allocation
We use five table datasets to find the trade-off between token allocation for table sampling and table augmentation, as demonstrated in Figure 3. We find that: (1) A balanced token distribution between the table and augmentation (around the 5:5 and 4:6 ratios, also known as a balanced T:A Ratio) generally yields the best performance for all five datasets. In indicates that properly controlling the token allocation can help LLMs achieve better performance. (2) Diminishing returns are observed when too many tokens are allocated to the augmentation information (as in the 3:7 ratio). This leads to a decrease in performance, suggesting that beyond a certain point, additional augmentation tokens may not be beneficial and could potentially detract from the main table content.
@@ 3.5 Trade-offs between Token Allocation
This trade-off highlights a broader concept in data processing and machine learning: the balance between information overload and information scarcity. Over-augmentation can lead to confusion and difficulty in discerning key patterns or insights. On the other hand, excessive sampling could result in an incomplete or biased understanding of the data. Note that the optimal T: A Ratio may vary across different datasets, as each may have unique characteristics making certain ratios more effective.
@@
Figure 4: Comparative Analysis of Model Performance Across TabFact, HybridQA, FEVEROUS and SQA. The series of graphs illustrates the frequency distribution of token lengths alongside the LLM performance (%) for three distinct methods: only sampling, only augmentation, and the hybrid method. Each subplot corresponds to a different dataset, depicting how table token length impacts model accuracy for various data augmentation and sampling strategies. Noted that “only augmentation” method refers to adding only the augmentation information to the prompt, without using any sampling method.
3.6 Large Table Analysis
Compared to the smaller-sized table, large table can grow to immense sizes, which make it more difficult to efficiently maintain, and reason over the tables. In designing TAP4LLM, performance optimization in this scenario is critical. We plot the distribution regarding the token numbers from the table across the five datasets in Figure 4, and also illustrate the impact of token numbers on LLM performance for three distinct settings.
@@ 3.6 Large Table Analysis
We can observe that: Shorter token lengths dominate the datasets, indicating a prevalence of text entries are relatively brief. Augmentation techniques excel with these shorter lengths by providing focused, enriched contexts that facilitate better model learning from simpler inputs. In contrast, sampling methods prove more effective for larger tables, suggesting they help manage data complexity by focusing on relevant data segments. The hybrid method, combining both techniques, shows consistent accuracy across various token lengths, highlighting its ability to leverage the strengths of both augmentation and sampling for improved performance across the board.
@@ 4 Related Work
Large Language Models for Tabular Data. Following the line of LLMs in natural language processing, researchers have also explored large models for various modalities like vision Gong et al. (2023); Kirillov et al. (2023) and speech Huang et al. (2023). From a technical standpoint, their ability to generate human-like text has opened new vistas of possibilities for processing tabular data. Nevertheless, it is non-trivial to directly employ the vanilla LLMs in the tabular area for two reasons: (i)-Global Table Understanding: the GPTs are known to suffer from the limited token length and thus, they can not read a whole large table, making them hard to understand the global tabular information. (ii)-Generalized to Tabular Domain: Second, their training processes are tailored for natural languages and thus, they are less generalizable when handling tabular data. There have been several works Hu et al. (2023); Zhong et al. (2017); Li et al. (2023b, a) developed to integrate natural language for tabular data analysis. Table Augmentation. Table augmentation is a technique used to improve the generalization performance and robustness of machine learning models. To enhance the performance and capabilities of LLMs in various domains, various explorations have been done to augment their knowledge grounding. It involves incorporating structured knowledge Sui et al. (2023); Xie et al. (2022), commonsense knowledge Bian et al. (2023); Ogundare and Araya (2023); Shen et al. (2023); Guo et al. (2023), and analytical knowledge He et al. (2023); Jena et al. (2022) into the pre-training and inference processes. For example, Jena et al. (2022) proposes to semi-automatically transform existing tabular data to create diverse/inventive natural language inference instances for better zero-shot performance. He et al. (2023) proposes a multi-tasking Metadata model that leverages field distribution and knowledge graph information to accurately infer analysis metadata for tables, and then demonstrates its deployment in a data analysis product for intelligent features. We follow the definition of statistical features from He et al. (2023). Each term with the corresponding definition is shown in Table 8.
@@ 5 Conclusion
We present TAP4LLM, a pre-processor designed to enhance the effectiveness of LLMs in tabular reasoning tasks. Technically, our method paves the way for interactive table reasoning as a natural language-driven framework, allowing for different components as plugins. Through three core components: table sampling, table augmentation, and table packing & serialization, we address several major challenges in comprehensive table understanding. We believe that TAP4LLM has the potential to enhance table modeling and exploratory data analysis (EDA) and enable various domains such as finance, transportation etc. Our study will be beneficial for table-based research and serve as an auxiliary tool to help better utilize LLMs on tabular-based/structured data-based tasks.
@@
Code generation-based methods Cheng et al. (2023); Gemmell and Dalton (2023) have been proposed to leverage LLMs to convert natural language queries into executable code or structured representations. We believe that semantic parsing or code generation is an important research direction. However, due to the page limits, we will leave this topic to further exploration. Additionally, our empirical study is mostly designed for English, rather than multilingual scenarios. The conversation on multilingual capabilities will also be part of future exploration.
@@
All the experiments in this paper run on GPU clusters with 8 NVIDIA A100 GPUs. Notably, all GPU clusters within our organization are shared instead of exclusive usage, and their carbon footprints are monitored in real-time. Our organization also consistently upgrade our data centers to reduce energy use.

### v3

@@ 2.2.2 Retrieval-based Augmentation
Large Language Models have occasionally been observed to generate hallucinated or factually inaccurate text Zhou et al. (2021); Zhao et al. (2023). To mitigate these issues, several works have proposed to strengthen LLMs with information retrieval systems Shi et al. (2023); Jiang et al. (2023b); Nakano et al. (2022), which enables LLMs to retrieve relevant content from an external repository (knowledge corpus). It has been verified that retrieval-augmented LLMs can generate texts in response to user input with fewer hallucinations Nakano et al. (2022). Furthermore, by incorporating customized private data resources, retrieval-augmented LLMs can respond to in-domain queries that cannot be answered by LLMs trained with public data. As previous works Nakano et al. (2022); Shi et al. (2023); Jiang et al. (2023b) suggested, LLMs can generate more factual answers by feeding the references retrieved from the external corpus.
In TAP4LLM, we have fortified the document retrieval capabilities of LLMs and consider two components: (1) document references: provide supplemental relevant web pages as the references for the given table; (2) term explanation: explain strange/ambiguous term in the given table. We utilize technologies including vector databases Wang et al. (2021a) and LangChain LangChain (2022) to facilitate the retrieval of pertinent information from Wikipedia [^1] https://www.wikipedia.org/. The details for document references and term explanation can be found in Appendix B.
@@ 2.2.3 Self-consistency-based Augmentation
We follow Sui et al. (2023) to implement the self-consistency-based augmentation approach. Specifically, we append the instruction "Identify critical values and ranges of the last table related to the statement" to the initial prompt and forward it to the LLM. The output generated from this instruction is then incorporated back into the prompt. Following this, we then re-forward the enriched prompt to LLMs — containing both the initial query and the newly generated insights, along with task-specific instructions for further processing.
@@ 2.3 Table Packing
The table packing module is motivated by the need to preserve efficient reasoning without altering the LLM architecture. First, we apply this module to manage token-limit allocation for the table sampling and augmentation. To achieve this, we conduct an empirical study to determine the optimal ratio between the sub-table length and the length of the augmentation information, as illustrated in Figure 3. The packing process is regulated by a user-defined token limit parameter, which sets the maximum truncated token length. Additionally, we are inspired by the study Sui et al. (2023), which highlights that using markup languages such as HTML or XML significantly improves generation quality in comparison to TQA and TFV. In line with this, TAP4LLM supports multiple serialization formats, including HTML, XML, JSON, CSV, NL+Sep (a common option, e.g., using ‘||||’ as a cell/column separator), and Markdown, etc.
Table 2: Comparative results of the table augmentation methods. We use a semantic-based sampling method without augmentation as the baseline. The term “Delta” refers to the performance gap between each method and the baseline. The top-3 performance gaps on each dataset are highlighted in green, with the best performance being underlined. Note that since only the ToTTo dataset contains hierarchical headers, we only provide the “header hierarchy” method on this dataset. “D/M + SF" refers to Dimension/Measure+ Semantic Field Type.
@@
3 Experiments
In this section, we first present the experimental setup, followed by an extensive comparison between baselines within each module in TAP4LLM. Additionally, we conduct an ablation study and provide a thorough evaluation of the performance of TAP4LLM. For further details on the experimental setup and additional experiments, please refer to Appendix D and E.
3.1 Experiment Settings
Datasets. We evaluate TAP4LLM on five TQA & TFV datasets: Sequential Question Answering (SQA) Iyyer et al. (2017), HybridQA Chen et al. (2020b), TabFact Chen et al. (2020a), ToTTo Parikh et al. (2020). Additionally, we set up TAP4LLM on a Text-to-SQL database Spider Yu et al. (2018), detailed in  E.4. The statistics of the datasets are given in Table 6, and the details of the datasets and metrics are described in Appendix D.1.
Models. We select state-of-the-art LLMs that have been widely studied in text generation and reasoning, including multiple GPT-series models and advanced open-source LLMs. Details of the tested models and embedding methods are provided in Appendix D.2, while the experimental results for open-source LLMs are available in Appendix E.3.
3.2 Results of Table Sampling
@@
As shown in Table 1, we perform comparative experiments on various table sampling methods, leading to the following key observations: (1) Semantic-based sampling with column grounding outperforms other sampling methods across all datasets by effectively selecting table parts most relevant to queries. Centroid-based sampling also shows competitive results by clustering data points within tables, though it lacks query-table relevance consideration. Moreover, when combining these two strong variants (hybrid sampling), it shows the most powerful capability. (2) The rule-based sampling method content snapshot, while not as precise in capturing query-specific information, offers a promising, efficient alternative by focusing on essential table content through $n$-gram overlap, without the need for complex embedding calculations. (3) Direct encoding methods, including using GPT-3.5-turbo with a 32k token limit or a 4k token-sized context window with truncation, demonstrate inferior performance. This suggests that while they can encompass more table information, they may introduce noise or lose critical context, undermining the table reasoning process and highlighting the importance of strategic sampling for optimal LLM performance.
3.3 Results of Table Augmentation
@@
For the comparative experiments of table augmentation methods, we use the semantic-based sampling method as the baseline and report the performance gap between adding each augmentation method or not. Table 2 provides several key insights, summarized as follows: (1) Table augmentation methods further improve LLM’s reasoning ability after sampling. For example, D/M + SF achieves higher accuracy across all six datasets (a most significant increase on TabFact +2.87%). Docs References and Term Explanations add meaningful context and semantic understanding to the model’s processing of tables, with (SQA +5.13%, ToTTo +4.07% ). The Self-Prompting further exemplifies the potential for iterative improvement in query and response generation. However, not all augmentation methods yield positive outcomes. Table Size offers minimal performance enhancement and Header Hierarchy shows that introducing a hierarchy may complicate the model’s ability to process the tabular information in some contexts, possibly by adding unnecessary complexity. (2) Additionally, the comparison of cell selection methods for Term Explanations highlights the superior performance of LLM-based selection over heuristic approaches. We find that LLM-based cell selection outperforms the heuristics-based cell selection with improvements in “Delta” ranging from 0.80% to 4.07%. While achieving higher performance, the LLM-based method also increases the calling budget as it requires additional LLM calls. These results indicate that the method’s effectiveness varies with the dataset. i.e. It’s beneficial for datasets requiring complex text understanding and generation (SQA and ToTTo). However, its impact is less distinct or even slightly negative in datasets involving different types of data or nuanced tasks (FEVEROUS and HybridQA).
@@
Through the experiment results, we also observe that different augmentation methods perform well on the same dataset. For example, D/M + SF, Statistics Feature, Term Explanation and Self Prompting all show significant improvement on the TabFact dataset. This suggests that combining multiple augmentation methods may have cumulative effects, leading to improved performance. In this work, we only report one simple intuitive approach - appending all augmentation information together into the prompt. We leave more fine-grained combinations for future exploration.
@@ 3.4 Ablation Study of TAP4LLM
As shown in Table 3, we conduct an ablation study to evaluate the impact of various components on the performance of TAP4LLM. Each row represents the model’s performance with the removal of a specific component. Our findings indicate that all components contribute to the overall effectiveness of the model, with certain components, such as table sampling and table augmentation, being particularly critical. The study also reveals that each dataset responds differently to the removal of features, underscoring the importance of a tailored design when optimizing for specific datasets. We also report the performance using the most optimal combination of table sampling and augmentation for each dataset, as presented in Table 3.
@@
3.5 Trade-offs between Token Allocation
We employ five table datasets to explore the trade-off between token allocation for table sampling and table augmentation, as illustrated in Figure 3. We observe that: (1) A balanced token distribution between the table and augmentation (approximately 5:5 or 4:6, referred to as the balanced T:A ratio) generally achieves the best performance across all five datasets. This suggests that carefully managing token allocation can enhance LLM performance. (2) Diminishing returns are observed when an excessive number of tokens are allocated to augmentation information (e.g., a 3:7 ratio), leading to a decline in performance. This indicates that beyond a certain point, additional augmentation tokens may no longer be beneficial and could detract from the core table content.
@@ 3.5 Trade-offs between Token Allocation
The trade-off we analyze above reflects a broader principle in data processing and machine learning: the balance between information overload and information scarcity. Over-augmentation can introduce noise, making it harder to identify key patterns or insights, while excessive sampling may lead to an incomplete or biased understanding of the data. It is important to note that the optimal T:A ratio may vary across datasets, as each has unique characteristics that make certain ratios more effective.
@@
Figure 4: Comparative Analysis of Model Performance Across TabFact, HybridQA, FEVEROUS and SQA. The series of graphs illustrates the frequency distribution of token lengths alongside the LLM performance (%) for three distinct methods: only sampling, only augmentation, and the hybrid method. Each subplot corresponds to a different dataset, depicting how table token length impacts model accuracy for various data augmentation and sampling strategies. Note that the “only augmentation” method refers to adding only the augmentation information to the prompt, without using any sampling method.
3.6 Computational Implications
Table 4 demonstrates the bottleneck of computation for TAP4LLM. Specifically, The first row calculates the average LLM calls conducted during the Table Augmentation with Docs References and Term Explanations, which contributes to $N+C$ times LLM calling. Here, $N$ is the number of terms detected by the LLM, and $C$ is a constant equal to 2 if there is no error occurs when calling the API. We also measure the token usage for each LLM call during the Table Sampling. Since the Table Sampling module only retrieves the top-k-related rows/columns to reconstruct a sub-table for LLM calling, this ensures token expenditure without compromising the model’s performance, especially in cases where thousands of rows exist in a table. The second row of the table 4 shows the average token usage per question needed by TAP4LLM after table sampling.
Table 4: Efficiency of multiple preprocessing steps in TAP4LLM
Dataset SQA FEVEROUS TabFact HybridQA
Average LLM Calls 4.7 5.2 5.3 8.4
Average Token Usage 637 512 417 742
3.7 Large Table Analysis in TAP4LLM
@@
Compared to smaller tables, large tables can grow to immense sizes, making them more difficult to maintain and reason over tabular data. In designing TAP4LLM, performance optimization in this context is essential. Figure 4 presents the distribution of token counts from the table across the five datasets, while also illustrating the impact of token numbers on LLM performance in three distinct settings. We can observe that: (1) Shorter token lengths dominate the datasets, indicating the prevalence of text entries is relatively brief. (2) Augmentation techniques excel with these shorter lengths by providing focused, enriched contexts that facilitate better model learning from simpler inputs. In contrast, sampling methods prove more effective for larger tables, suggesting that they help manage data complexity by focusing on relevant data segments. (2) The hybrid method shows consistent performance across various token lengths, highlighting its ability to leverage the strengths of both augmentation and sampling for robust performance enhancement.
@@ 4 Related Work
Large Language Models for Tabular Data
Following the line of LLMs in natural language processing, researchers have also explored large models for various modalities like vision Gong et al. (2023); Kirillov et al. (2023) and speech Huang et al. (2023). From a technical standpoint, their ability to generate human-like text has opened new vistas of possibilities for processing tabular data. Nevertheless, it is non-trivial to directly employ the vanilla LLMs in the tabular area for two reasons: (i)-Global Table Understanding: the GPTs are known to suffer from the limited token length and thus, they can not read a whole large table, making them hard to understand the global tabular information. (ii)-Generalized to Tabular Domain: Second, their training processes are tailored for natural languages and thus, they are less generalizable when handling tabular data. There have been several works Hu et al. (2023); Zhong et al. (2017); Li et al. (2023b, a) developed to integrate natural language for tabular data analysis.
Table Augmentation
Table augmentation is a technique used to improve the generalization performance and robustness of machine learning models. To enhance the performance and capabilities of LLMs in various domains, various explorations have been done to augment their knowledge grounding. It involves incorporating structured knowledge Sui et al. (2023); Xie et al. (2022), commonsense knowledge Bian et al. (2023); Ogundare and Araya (2023); Shen et al. (2023); Guo et al. (2023), and analytical knowledge He et al. (2023); Jena et al. (2022) into the pre-training and inference processes. For example, Jena et al. (2022) proposes to semi-automatically transform existing tabular data to create diverse/inventive natural language inference instances for better zero-shot performance. He et al. (2023) proposes a multi-tasking Metadata model that leverages field distribution and knowledge graph information to accurately infer analysis metadata for tables, and then demonstrates its deployment in a data analysis product for intelligent features.
@@ 5 Conclusion
In this paper, we propose TAP4LLM (Table Provider for LLM) as a powerful toolkit designed to enhance the interaction between LLMs and structured table data. It provides optimized prompt designs and robust functionalities to ensure high-quality outputs when LLMs process table-related inputs. We believe that TAP4LLM has the potential to significantly improve table modeling and exploratory data analysis (EDA), with applications in various domains.
@@
Code generation-based methods Cheng et al. (2023); Gemmell and Dalton (2023); He et al. (2024) have been proposed to leverage LLMs to convert natural language queries into executable code or structured representations. We believe that semantic parsing or code generation is an important research direction. However, due to the page limits, we will leave this topic to further exploration. Additionally, our empirical study is mostly designed for English, rather than multilingual scenarios. The conversation on multilingual capabilities will also be part of future exploration.
@@
All the datasets used in this paper are public and have been reviewed to ensure they do not contain any personally identifiable information or offensive content. However, as these datasets are sourced from the Internet, potential bias may still be present. Furthermore, despite our careful review, the process of table augmentation with LLMs throughout may inadvertently introduce inappropriate information into the preprocessed data. In addition, all the experiments in this paper are run on GPU clusters with 8 NVIDIA A100 GPUs.